# ReSplat: Learning Recurrent Gaussian Splats

### v1

@@ Abstract
While feed-forward Gaussian splatting models provide computational efficiency and effectively handle sparse input settings, their performance is fundamentally limited by the reliance on a single forward pass during inference. We propose ReSplat, a feed-forward recurrent Gaussian splatting model that iteratively refines 3D Gaussians without explicitly computing gradients. Our key insight is that the Gaussian splatting rendering error serves as a rich feedback signal, guiding the recurrent network to learn effective Gaussian updates. This feedback signal naturally adapts to unseen data distributions at test time, enabling robust generalization. To initialize the recurrent process, we introduce a compact reconstruction model that operates in a $16\times$ subsampled space, producing $16\times$ fewer Gaussians than previous per-pixel Gaussian models. This substantially reduces computational overhead and allows for efficient Gaussian updates. Extensive experiments across varying of input views (2, 8, 16), resolutions ($256\times 256$ to $540\times 960$), and datasets (DL3DV and RealEstate10K) demonstrate that our method achieves state-of-the-art performance while significantly reducing the number of Gaussians and improving the rendering speed. Our project page is at haofeixu.github.io/resplat.
Figure 1: Learning recurrent Gaussian splats in a feed-forward manner. We propose ReSplat, a feed-forward recurrent network that iteratively refines 3D Gaussian splats to improve sparse view settings where optimization-based 3DGS (Kerbl et al., 2023) struggles. As initialization (iteration 0), we introduce a compact reconstruction model that predicts Gaussians in a $16\times$ subsampled space, producing $16\times$ fewer Gaussians and $4\times$ faster rendering than per-pixel MVSplat (Chen et al., 2024b) and DepthSplat (Xu et al., 2025). The reduced number of Gaussians makes subsequent refinement efficient. Compared to the optimization-based 3DGS, ReSplat is $100\times$ faster thanks to its feed-forward design, while still benefiting from iterative updates. Here we show results for 8 input views ($512\times 960$ resolution) on DL3DV (Ling et al., 2023) dataset; see Table˜1 for detailed metrics.
@@ 1 Introduction
Based on this compact initial reconstruction, we train a weight-sharing recurrent network that iteratively improves the reconstruction. Crucially, the recurrent network leverages the rendering error of input views to determine where and how to update the Gaussians. Specifically, we render the input views (available at test time) using the current prediction, compute the rendering error, and propagate it in feature space to 3D Gaussians. The recurrent network then predicts the parameter updates directly from this error and the current Gaussians, without requiring explicit gradient computation.
We validate our method through extensive experiments across diverse scenarios. On the challenging DL3DV (Ling et al., 2023) dataset, using 8 input views at $512\times 960$ resolution, our learned recurrent model improves PSNR by +2.7 dB, while using only $1/16$ of the Gaussians and achieving $4\times$ faster rendering speed. We also demonstrate that our recurrent model leads to robust generalization to unseen datasets and image resolutions, where previous single-step feed-forward models usually struggle. With 16 input views at $540\times 960$ resolution, we outperform Long-LRM (Chen et al., 2025b) while using $4\times$ fewer Gaussians. On the commonly used two-view RealEstate10K benchmark, our ReSplat also achieves state-of-the-art results, demonstrating the strong performance of our method.
@@ 2 Related Work
Figure 2: Learning to recurrently update 3D Gaussians. Given NN posed input images, we first predict per-view depth maps at $1/4$ resolution and then unproject and transform them to a point cloud with image features $\{({\bm{p}}_{j},{{\bm{f}}}_{j})\}_{j=1}^{M}$, where $M=N\times\frac{HW}{16}$ is the number of points. We then reconstruct an initial set of 3D Gaussians $\{({\bm{g}}^{0}_{j},{\bm{z}}^{0}_{j})\}_{j=1}^{M}$ in a $16\times$ subsampled 3D space with a kNN and global attention-based Gaussian regressor. Next, we learn to refine the initial Gaussians recurrently. At each recurrent step $t$, we use the current Gaussian prediction to render input views and then compute the feature-level rendering errors $\bm{E}$ between rendered and ground-truth input views. A global attention is next applied on the rendering error to propagate the rendering errors to the 3D Gaussians. A kNN attention-based update module next takes as input the concatenation of current Gaussian parameters ${\bm{g}}^{t}_{j}$, the hidden state ${\bm{z}}^{t}_{j}$, and the rendering error ${\bm{e}}^{t}_{j}$, and predicts the incremental updates $\Delta{\bm{g}}^{t}_{j}$ and $\Delta{\bm{z}}^{t}_{j}$. We iterate this process until a total number of $T$ steps.
@@ 3.1 Initial Gaussian Reconstruction
Subsampled 3D Space. Our initial Gaussian reconstruction model is based on the DepthSplat (Xu et al., 2025) architecture, which includes both multi-view and monocular branches for per-pixel Gaussian prediction. However, unlike DepthSplat, we predict Gaussians in a spatially $16\times$ subsampled 3D space ($N\times\frac{HW}{16}$), and thus our number of Gaussians is $16\times$ fewer than DepthSplat. To achieve $16\times$ subsampling, we resize the full resolution depth predictions from the depth model in DepthSplat to $1/4$ resolution ($N\times\frac{H}{4}\times\frac{W}{4}$), and then unproject and transform them in 3D via camera parameters to obtain a point cloud with $M=N\times\frac{HW}{16}$ points. Each 3D point ${\bm{p}}_{j}\in\mathbb{R}^{3}$ is also associated with a feature vector ${\bm{f}}_{j}\in\mathbb{R}^{C_{1}}$ extracted from the input images. This process can be expressed as
@@
Decoding to Gaussians. We use the point cloud $\{{\bm{p}}_{j}\}_{j=1}^{M}$ as the Gaussian centers and other Gaussian parameters are decoded from the 3D context-aggregated features $\{{\bm{f}}^{*}_{j}\}_{j=1}^{M}$. Accordingly, we obtain an initial set of 3D Gaussians with parameters $\{(\bm{\mu}_{j},\alpha_{j},\bm{\Sigma}_{j},\bm{\mathrm{sh}}_{j})\}_{j=1}^{M}$ and feature vectors $\{{\bm{f}}^{*}_{j}\}_{j=1}^{M}$. We use ${\bm{g}}^{0}_{j}\in\mathbb{R}^{C_{2}}$ to denote the concatenation of all the Gaussian parameters $(\bm{\mu}_{j},\alpha_{j},\bm{\Sigma}_{j},\bm{\mathrm{sh}}_{j})$ for the $j$-th Gaussian at initialization, where $C_{2}$ is the total number of parameters for each Gaussian. We use ${\bm{z}}^{0}_{j}$ to denote the initial hidden state of the $j$-th Gaussian for the subsequent recurrent process, and initialize it with feature ${\bm{f}}^{*}_{j}$: ${\bm{z}}^{0}_{j}={\bm{f}}^{*}_{j}\in\mathbb{R}^{C_{1}}$. Thus, the initial Gaussians can be represented as
@@ 3.2 Recurrent Gaussian Update
Computing the Rendering Error. Given that we have access to the input views at test time, we are able to create a feedback loop to guide the recurrent network to learn the incremental updates. More specifically, we first render the input views $\{\hat{\bm{I}}^{t}_{i}\}_{i=1}^{N}$ based on the current Gaussian parameters at iteration $t$, and then measure the difference between the rendered and ground-truth input views. We observe that measuring the rendering error in the feature space performs better than in the raw RGB space. In particular, we extract the first-three-stage features (at $1/2$, $1/4$ and $1/8$ resolutions) of the ImageNet (Deng et al., 2009) pre-trained ResNet-18 (He et al., 2016) for the rendered input views and ground-truth input views, and bilinearly resize the three-scale features to the same $1/4$ resolution, followed by concatenation. We denote the extracted features as $\{\hat{\bm{F}}^{t}_{i}\}_{i=1}^{N}$ and $\{{\bm{F}}_{i}\}_{i=1}^{N}$ ($\hat{\bm{F}}^{t}_{i},{\bm{F}}_{i}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times C_{3}}$) for rendered and ground-truth input views, respectively. We then compute the difference between features with subtraction $\{\hat{\bm{F}}^{t}_{i}-{\bm{F}}_{i}\}_{i=1}^{N}$. We denote all rendering errors as $\hat{\bm{E}}^{t}=\{\hat{\bm{e}}^{t}_{j}\}_{j=1}^{N\times\frac{H}{4}\times\frac{W}{4}}$, where $\hat{\bm{e}}^{t}_{j}\in\mathbb{R}^{C_{3}}$ is the $j$-th feature difference of dimension $C_{3}$ at iteration $t$.
Propagating the Rendering Error to Gaussians. To propagate the rendering error to 3D Gaussians such that they can guide the network to update the Gaussians, a straightforward approach is to concatenate the rendering error $\hat{\bm{e}}^{t}_{j}$ with the Gaussians (${\bm{g}}^{t}_{j},{\bm{z}}^{t}_{j}$) in a spatially aligned manner since they have the same number of points ($N\times\frac{HW}{16}$). However, in this way, the $j$-th Gaussian can only receive information about the $j$-th rendered pixel, while it can also contribute to other rendered pixels during the rendering process. To propagate the rendering error more effectively, we propose to apply global attention on all the $N\times\frac{H}{4}\times\frac{W}{4}$ rendering errors $\hat{\bm{E}}^{t}$, which enables each Gaussian to receive information from all rendering errors. This process can be formulated as
${\bm{E}}=\mathrm{global\_attention}(\hat{\bm{E}})=\{{\bm{e}}^{t}_{j}\}_{j=1}^{N\times\frac{HW}{16}},$ (5)
@@
Recurrent Gaussian Update. Let the Gaussians at iteration $t$ be $\mathcal{G}^{t}=\{({\bm{g}}^{t}_{j},{\bm{z}}^{t}_{j})\}_{j=1}^{M}$, our update module predicts the incremental updates of Gaussian parameters and hidden state as
@@
and then they are added to the current prediction (Eq.˜4). This process is iterated for $T$ times. We observe that our model converges after 3 iterations. During training, we randomly sample the number of iterations $T$ between 1 and 3, and our model supports different number of iterations at inference time, allowing a flexible trade-off between accuracy and speed with a single model. Since the recurrent process occurs in the 3D space, we choose to use four kNN attention (Zhao et al., 2021) blocks as the recurrent architecture to model the local structural details.
@@ 4 Experiments
Implementation Details. We implement our method in PyTorch and use Flash Attention 3 (Shah et al., 2024) for efficient attention computations. We choose $k=16$ for kNN attentions following Point Transformer (Zhao et al., 2021), and our Gaussian splatting renderer is based on gsplat (Ye et al., 2025b)’s Mip-Splatting (Yu et al., 2024) implementation. We optimize our model with AdamW (Loshchilov & Hutter, 2017) optimizer with cosine learning rate schedule.
@@
Our default model uses a ViT-B (Dosovitskiy et al., 2020; Yang et al., 2024) backbone as the monocular branch of our depth prediction model (Xu et al., 2025), which has 223M parameters in total (209M for the initialization model and 14M for the recurrent model). For ablation experiments, we use a ViT-S backbone to save compute, which has 76M parameters in total (62M for the initialization model and 14M for the recurrent model).
@@
Evaluation Setup. To compare with previous methods, we mainly consider three evaluation setups. First, view synthesis from 8 input views at $512\times 960$ resolution on DL3DV, where we re-train 3DGS (Kerbl et al., 2023), MVSplat (Chen et al., 2024b) and DepthSplat (Xu et al., 2025) with their public code for fair comparisons. Second, view synthesis from 16 input views at $540\times 960$ resolution on DL3DV following Long-LRM (Chen et al., 2025b), with which we could perform comparisons with Long-LRM since its model weights are not released and re-training would be expensive. Third, we also evaluate on the commonly used 2-view ($256\times 256$) setup on RealEstate10K, where we compare with related methods like GS-LRM (Zhang et al., 2024a) and LVSM (Jin et al., 2025).
Training Details. For experiments on DL3DV, we adopt a progressive training strategy with gradually increased image resolutions and number of input views for better efficiency. More specifically, we first train our model with 8 input views at $256\times 448$ resolution, and then we fine-tune the model with 8 input views at $512\times 960$ resolution, and finally we fine-tune the model with 16 input views at $512\times 960$ resolution. For each stage, we train with 16 GH200 GPUs for 80K steps, with 50K steps for the initial reconstruction model and 30K steps for the recurrent model. For experiments on RealEstate10K at $256\times 256$ resolution, we first train the initial model with 16 GH200 GPUs for 200K steps and then train the recurrent model for 100K steps.

### v2

@@ Abstract
While feed-forward Gaussian splatting models offer computational efficiency and can generalize to sparse input settings, their performance is fundamentally constrained by relying on a single forward pass for inference. We propose ReSplat, a feed-forward recurrent Gaussian splatting model that iteratively refines 3D Gaussians without explicitly computing gradients. Our key insight is that the Gaussian splatting rendering error serves as a rich feedback signal, guiding the recurrent network to learn effective Gaussian updates. This feedback signal naturally adapts to unseen data distributions at test time, enabling robust generalization across datasets, view counts and image resolutions. To initialize the recurrent process, we introduce a compact reconstruction model that operates in a $16\times$ subsampled space, producing $16\times$ fewer Gaussians than previous per-pixel Gaussian models. This substantially reduces computational overhead and allows for efficient Gaussian updates. Extensive experiments across varying of input views (2, 8, 16, 32), resolutions ($256\times 256$ to $540\times 960$), and datasets (DL3DV, RealEstate10K and ACID) demonstrate that our method achieves state-of-the-art performance while significantly reducing the number of Gaussians and improving the rendering speed. Our project page is at haofeixu.github.io/resplat.
Figure 1: Learning recurrent Gaussian splats in a feed-forward manner. We propose ReSplat, a feed-forward recurrent network that iteratively refines 3D Gaussian splats to improve sparse view settings where optimization-based 3DGS (Kerbl et al., 2023) struggles. As initialization (iteration 0), we introduce a compact reconstruction model that predicts Gaussians in a $16\times$ subsampled space, producing $16\times$ fewer Gaussians and $4\times$ faster rendering than per-pixel MVSplat (Chen et al., 2024b) and DepthSplat (Xu et al., 2025). The reduced number of Gaussians makes subsequent refinement efficient. Compared to the optimization-based 3DGS, ReSplat is $100\times$ faster thanks to its feed-forward design, while still benefiting from iterative updates. Here we show results for 8 input views ($512\times 960$ resolution) on DL3DV dataset; see Table˜1 for detailed metrics.
@@ 1 Introduction
Based on this compact initial reconstruction, we train a weight-sharing recurrent network that iteratively improves the reconstruction. Crucially, the recurrent network leverages the rendering error of input views to determine where and how to update the Gaussians. Specifically, we render the input views (available at test time) using the current prediction, compute the rendering error, and propagate it to 3D Gaussians. The recurrent network then predicts the parameter updates directly from this error and the current Gaussians, without requiring explicit gradient computation.
We validate our method through extensive experiments across diverse scenarios. On the challenging DL3DV (Ling et al., 2023) dataset, using 8 input views at $512\times 960$ resolution, our learned recurrent model improves PSNR by +3.5dB, while using only $1/16$ of the Gaussians and achieving $4\times$ faster rendering speed. We also demonstrate that our recurrent model leads to robust generalization to unseen datasets, view counts, and image resolutions, where previous single-step feed-forward models usually struggle. With 16 input views at $540\times 960$ resolution, we outperform Long-LRM (Chen et al., 2025b) by +0.8dB PSNR while using $4\times$ fewer Gaussians. On the commonly used two-view RealEstate10K (Zhou et al., 2018) and ACID (Liu et al., 2021) benchmarks, our ReSplat also achieves state-of-the-art results, demonstrating the strong performance of our method.
@@ 2 Related Work
Figure 2: Learning to recurrently update 3D Gaussians. Given NN posed input images, we first predict per-view depth maps at $1/4$ resolution and then unproject and transform them to a point cloud with image features $\{({\bm{p}}_{j},{{\bm{f}}}_{j})\}_{j=1}^{M}$, where $M=N\times\frac{HW}{16}$ is the number of points. We then reconstruct an initial set of 3D Gaussians $\{({\bm{g}}^{0}_{j},{\bm{z}}^{0}_{j})\}_{j=1}^{M}$ in a $16\times$ subsampled 3D space with a kNN and global attention-based Gaussian regressor. Next, we learn to refine the initial Gaussians recurrently. At each recurrent step $t$, we use the current Gaussian prediction to render input views and then compute the rendering errors $\hat{\bm{E}}^{t}$ between rendered and ground-truth input views. A global attention is next applied on the rendering error to propagate the rendering errors to the 3D Gaussians. A kNN attention-based update module next takes as input the concatenation of current Gaussian parameters ${\bm{g}}^{t}_{j}$, the hidden state ${\bm{z}}^{t}_{j}$, and the rendering error ${\bm{e}}^{t}_{j}$, and predicts the incremental updates $\Delta{\bm{g}}^{t}_{j}$ and $\Delta{\bm{z}}^{t}_{j}$. We iterate this process until a total number of $T$ steps.
@@ 3.1 Initial Gaussian Reconstruction
Subsampled 3D Space. Our initial Gaussian reconstruction model is based on the DepthSplat (Xu et al., 2025) architecture. However, unlike DepthSplat, we predict Gaussians in a spatially $16\times$ subsampled 3D space ($N\times\frac{HW}{16}$), and thus our number of Gaussians is $16\times$ fewer than DepthSplat. To achieve $16\times$ subsampling, we resize the full resolution depth predictions from the depth model in DepthSplat to $1/4$ resolution ($N\times\frac{H}{4}\times\frac{W}{4}$), and then unproject and transform them in 3D via camera parameters to obtain a point cloud with $M=N\times\frac{HW}{16}$ points. Each 3D point ${\bm{p}}_{j}\in\mathbb{R}^{3}$ is also associated with a feature vector ${\bm{f}}_{j}\in\mathbb{R}^{C_{1}}$ extracted from the input images:
@@
Decoding to Gaussians. We use the point cloud $\{{\bm{p}}_{j}\}_{j=1}^{M}$ as the Gaussian centers and other Gaussian parameters are decoded with a lightweight Gaussian head (two-layer MLP) from the 3D context-aggregated features $\{{\bm{f}}^{*}_{j}\}_{j=1}^{M}$. Accordingly, we obtain an initial set of 3D Gaussians with parameters $\{(\bm{\mu}_{j},\alpha_{j},\bm{\Sigma}_{j},\bm{\mathrm{sh}}_{j})\}_{j=1}^{M}$ and feature vectors $\{{\bm{f}}^{*}_{j}\}_{j=1}^{M}$. We use ${\bm{g}}^{0}_{j}\in\mathbb{R}^{C_{2}}$ to denote the concatenation of all the Gaussian parameters $(\bm{\mu}_{j},\alpha_{j},\bm{\Sigma}_{j},\bm{\mathrm{sh}}_{j})$ for the $j$-th Gaussian at initialization, where $C_{2}$ is the total number of parameters for each Gaussian. We use ${\bm{z}}^{0}_{j}$ to denote the initial hidden state of the $j$-th Gaussian for the subsequent recurrent process, and initialize it with feature ${\bm{f}}^{*}_{j}$: ${\bm{z}}^{0}_{j}={\bm{f}}^{*}_{j}\in\mathbb{R}^{C_{1}}$. Thus, the initial Gaussians can be represented as
@@ 3.2 Recurrent Gaussian Update
Computing the Rendering Error. Given that we have access to the input views at test time, we are able to create a feedback loop to guide the recurrent network to learn the incremental updates. More specifically, we first render the input views $\{\hat{\bm{I}}^{t}_{i}\}_{i=1}^{N}$ based on the current Gaussian parameters at iteration $t$, and then measure the difference between the rendered and ground-truth input views. We evaluate several different methods to compute the rendering error and observe that a combination of pixel-space and feature-space rendering errors performs best.
In particular, we first use $\{\hat{\bm{I}}^{t}_{i}-{\bm{I}}_{i}\}_{i=1}^{N}$ to measure the rendering error in the pixel space, and then perform $4\times$ spatial downsampling with pixel unshuffle to align with the number of 3D Gaussians. For the feature-space rendering error, we extract the first-three-stage features (at $1/2$, $1/4$ and $1/8$ resolutions) of the ImageNet (Deng et al., 2009) pre-trained ResNet-18 (He et al., 2016) for the rendered input views and ground-truth input views, and bilinearly resize the three-scale features to the same $1/4$ resolution, followed by concatenation. We denote the extracted features as $\{\hat{\bm{F}}^{t}_{i}\}_{i=1}^{N}$ and $\{{\bm{F}}_{i}\}_{i=1}^{N}$ ($\hat{\bm{F}}^{t}_{i},{\bm{F}}_{i}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times C_{3}}$) for rendered and ground-truth input views, respectively. We then compute the difference between features with subtraction $\{\hat{\bm{F}}^{t}_{i}-{\bm{F}}_{i}\}_{i=1}^{N}$. We combine pixel-space and feature-space rendering errors via element-wise addition. To match channel dimensions, the pixel-space error is first projected to the feature space using a linear layer followed by LayerNorm (Ba et al., 2016). This process can be expressed as
$\hat{\bm{E}}^{t}=f\left(\{\hat{\bm{I}}_{i}^{t}\}_{i=1}^{N},\{\bm{I}_{i}\}_{i=1}^{N}\right)=\{\hat{\bm{F}}^{t}_{i}-{\bm{F}}_{i}\}_{i=1}^{N}+\text{proj}(\{\hat{\bm{I}}^{t}_{i}-{\bm{I}}_{i}\}_{i=1}^{N}),$ (5)
where “proj” is the operation mentioned before to match dimensions. We denote all rendering errors as $\hat{\bm{E}}^{t}=\{\hat{\bm{e}}^{t}_{j}\}_{j=1}^{N\times\frac{H}{4}\times\frac{W}{4}}$, where $\hat{\bm{e}}^{t}_{j}\in\mathbb{R}^{C_{3}}$ is the $j$-th feature difference of dimension $C_{3}$ at iteration $t$.
Propagating the Rendering Error to Gaussians. To propagate the rendering error to 3D Gaussians such that they can guide the network to update the Gaussians. A straightforward approach is to concatenate the rendering error $\hat{\bm{e}}^{t}_{j}$ with the Gaussians (${\bm{g}}^{t}_{j},{\bm{z}}^{t}_{j}$) in a spatially aligned manner since they have the same number of points ($N\times\frac{HW}{16}$). However, in this way, the $j$-th Gaussian can only receive local information around the $j$-th rendered pixel, while it can also contribute to other rendered pixels during the rendering process. To propagate the rendering error more effectively, we propose to apply global attention on all the $N\times\frac{H}{4}\times\frac{W}{4}$ rendering errors $\hat{\bm{E}}^{t}$, which enables each Gaussian to receive information from all rendering errors. This process can be formulated as:
${\bm{E}}^{t}=\mathrm{global\_attention}(\hat{\bm{E}}^{t})=\{{\bm{e}}^{t}_{j}\}_{j=1}^{N\times\frac{HW}{16}},$ (6)
@@
Recurrent Gaussian Update. Let the Gaussians at iteration $t$ be $\mathcal{G}^{t}=\{({\bm{g}}^{t}_{j},{\bm{z}}^{t}_{j})\}_{j=1}^{M}$, our update module predicts the incremental updates of Gaussian parameters and hidden state as:
@@
and then they are added to the current prediction (Eq.˜4). This process is iterated $T$ times. We observe that our model converges after 4 iterations. During training, we randomly sample the number of iterations $T$ between 1 and 4, and our model supports different number of iterations at inference time, allowing a flexible trade-off between accuracy and speed with a single model. Since the recurrent process occurs in the 3D space, we choose to use four kNN attention (Zhao et al., 2021) blocks as the recurrent architecture to model the local structural details. The Gaussian updates ${\bm{g}}^{t}_{j}$ are decoded with a lightweight update head (four-layer MLP).
@@ 4 Experiments
Implementation Details. We implement our method in PyTorch and use Flash Attention 3 (Shah et al., 2024) for efficient attention computations. We choose $k=16$ for kNN attentions following Point Transformer (Zhao et al., 2021), and our Gaussian splatting renderer is based on gsplat (Ye et al., 2025b)’s Mip-Splatting (Yu et al., 2024) implementation. We optimize our model with AdamW (Loshchilov & Hutter, 2017) optimizer. More training details are presented in the appendix.
@@
Our default model uses a ViT-B (Dosovitskiy et al., 2020) backbone as part of our depth prediction model, which has 223M parameters in total (209M for the initialization model and 14M for the recurrent model). For ablation experiments, we use a ViT-S backbone to save compute, which has 76M parameters in total (62M for the initialization model and 14M for the recurrent model).
@@
Coordinate System. Since our recurrent network operates within a global 3D space, the selection of a coordinate system is critical, as it directly determines the spatial distribution of the Gaussian’s centers. Our datasets consist of video sequences with camera poses estimated from COLMAP (Schonberger & Frahm, 2016). We evaluated aligning the global reference frame to the first, middle, and last views of the input images. Empirically, we observed that using the middle input view as the reference coordinate system yields the best performance (see LABEL:tab:ablation_coord_system). We posit that this centers the coordinate system, reducing the maximum transformation distance to the first and last input views and effectively balancing the spatial positions of the 3D Gaussians.
Evaluation Setup. We mainly consider three evaluation setups. First, view synthesis from 8 input views at $512\times 960$ resolution on DL3DV, where we re-train 3DGS (Kerbl et al., 2023), MVSplat (Chen et al., 2024b) and DepthSplat (Xu et al., 2025) with their public code for fair comparisons. Second, view synthesis from 16 input views at $540\times 960$ resolution on DL3DV following Long-LRM (Chen et al., 2025b), with which we could perform comparisons with Long-LRM since its model weights are not released and re-training would be expensive. Third, we also evaluate on the commonly used 2-view ($256\times 256$) setup on RealEstate10K and ACID, where we compare with related methods like GS-LRM (Zhang et al., 2024a) and LVSM (Jin et al., 2025).